# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table 14. Monitoring challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Performance Monitoring Consistency and Variability
S01 Performance Monitoring Ensuring sustained performance monitoring in dynamically varying cloud environments Performance Variation Monitoring Monitoring performance fluctuations in containerized applications through benchmarking periods
S02 Benchmarking Performance Variation and Cost Efficiency Monitoring Efficiently tracking performance disparities and cost-effectiveness for refined monitoring in benchmarking SDBO Optimization Maximizing user budget utility with the optimization feature in SDBO
S29 Website Performance and Availability Monitoring Implementing continuous oversight over website availability and performance Kubernetes Monitoring Utilizing Kubernetes for monitoring performance of containerized services in multi-cloud environment
S38 User Experience Variability Monitoring Addressing monitoring fluctuations in user experience based on diverse factors CSP QoS Metrics Collection Collecting Cloud Service Provider (CSP) Quality of Service (QoS) metrics for benchmark data in multi-cloud environment
S43 Independent Performance Metrics Monitoring Establishing application-independent performance metrics for Round Trip Time (RTT) and payload throughput Performance Evaluations Conducting performance evaluations and autonomic resource provisioning for multi-cloud and MEC platforms using containerized applications
S50 Comprehensive Network Performance Measurement Monitoring Ensuring comprehensive performance measurements monitoring and accounting for transmission delays in the network Server Selection Scheme Applying a server selection scheme to minimize latency in multi-cloud environment based on end-to-end performance measurements
S62 Microservice Performance Degradation Monitoring Addressing unpredictable performance degradation in containerized microservices Tail Latency Prediction Implementing a modeling approach to predict end-to-end tail latency in multi-layer cloud data of containerized applications
Category 2: Resource and Infrastructure Management Monitoring
S05 Network Communication Monitoring Managing service-to-service communications over a network Service Mesh Implementation Deploying a network service mesh for service-discovery, secure connectivity, and visibility in multi-cloud platforms
S09 System Image and Configuration Monitoring Monitoring of system images, remote software packages, usage history, and container execution, configuration, and monitoring. Docker Server Deployment Deploying containerized services using Docker Server API version 1.23 in multi-cloud platforms
S21 Vcluster and Framework Management Monitoring Monitoring and managing the performance and resource usage of each Vcluster and application framework can be complex when dealing with multiple tenants and different types of applications. LXC Implementation Using Lightweight Containers (LXC) for efficient resource allocation and scalability in multi-cloud environment
S24 Multi-Cloud Resource Utilization Monitoring Monitoring container performance and resource usage across multiple clouds Prometheus and Grafana Monitoring Implementing Prometheus and Grafana for monitoring performance and availability of containerized applications in multi-cloud environment
S51 Dynamic Resource Management Monitoring Applications should add and release resources dynamically by measuring and monitoring usage Aneka Platform Utilizing the Aneka platform for monitoring resources and managing dynamic resource acquisition and release in multi-cloud environment
Category 3: Security, Compliance, and Trust Management Monitoring
S04 Provider Host Information Access Monitoring Overcoming the restriction in accessing provider host information ML Monitoring Using machine learning to supervise performance of containerized applications and OS usage
S34 Continuous Security Control Monitoring Strengthening continuous monitoring and control over security MUSA Platform Applying the MUSA Security Assurance Platform for continuous monitoring and dynamic security enforcement of containerized applications
S52 Multi-Cloud Security Compliance Monitoring Monitoring security compliance consistently across multiple cloud providers MUSA Security Compliance Using the MUSA Security Assurance Platform to monitor multi-cloud application compliance and ensure security
S53 Service Continuity and Data Privacy Monitoring Ensuring data privacy, uninterrupted service, fault tolerance, and availability NoMISHAP Implementation Implementing the NoMISHAP solution to address monitoring challenges in multi-cloud environment with containerized applications
S83 Trust Relationship and Network Contact Monitoring Building trust relationships, collecting judgments, and maintaining networks of contacts Trust Bootstrapping Implementing a trust bootstrapping mechanism that combines endorsement and decision tree classification for containerized applications in multi-cloud platforms
Category 4: Application Optimization and Adaptability Monitoring
S20 Adaptive System Design and State Monitoring Implementing continuous design with dynamic adaptability; Enhancing system state evaluation through robust monitoring Cloud Modelling Framework Utilizing a Cloud Modelling Framework for continuous design and dynamic adjustment of multi-cloud containerized applications
S36 Partial Download Execution Monitoring Ensuring accurate execution of applications for partially downloaded files FogDocker Implementation Using FogDocker with low runtime overhead to ensure execution correctness of containerized applications
S44 Event-Driven Application Behavior Monitoring Modeling event-triggered behavior of application components TOSCA Event Modeling Applying TOSCA modeling for event sources, events, and functions in containerized applications
S68 Microservice Optimization Monitoring Enhancing surveillance over microservices across multiple cloud environments for optimization MiCADO Optimization Utilizing MiCADO for runtime monitoring and optimization of cloud applications in multi-cloud environment
S85 Real-Time Application Monitoring Real-time monitoring of applications for immediate corrective action DRA Framework Implementing the DRA framework with the Monitoring and Tracking Platform (M4) for containerized applications in multi-cloud platforms
Category 5: System Complexity and Standardization Monitoring
S10 Proprietary Monitoring Solution Standardization Overcoming the constraints of complex, proprietary monitoring solutions and fostering standardization Pervasive Monitoring Implementing pervasive monitoring solutions for containerized applications within multi-cloud PaaS
S25 Kernel Resource Isolation Mechanism Monitoring Addressing the limitations of resource isolation mechanisms in the Linux kernel Leakage Defense Identifying in-container leakage channels and implementing a two-stage defensive approach in multi-cloud environment
S27 Multi-Cloud Application Monitoring Complexity Simplifying monitoring processes in multi-cloud applications DECIDE DevOps Expansion Expanding the DECIDE DevOps framework for native multi-cloud containerized applications
S28 Monitoring Agent Coordination Enhancing the coordination of monitoring agents for improved system behavior Autonomic Management Implementing an autonomic management system with monitoring agents, a performance model, and adaptive actions for containerized applications
S33 Heterogeneous Fog Resource Management Monitoring Efficient management of diverse fog resources Kubernetes Utility Architecture Implementing a master-worker nodes architecture with Kubernetes Utility (Kubeadm) in multi-cloud environment
S42 Application Topology and Dependency Monitoring Enhancing coordinated adaptation through thorough understanding of application topologies and dependencies MicroCloud TOSCA Library Utilizing the MicroCloud TOSCA library for standardized description of topologies in multi-cloud containerized applications
S45 Elastic Cloud Service Monitoring Research Investigating elastic cloud service monitoring as an open research problem JCatascopia Monitoring Utilizing JCatascopia for multi-layer, platform-independent monitoring of cloud containerized applications
S46 Microservice Monitoring Diversity Addressing diverse monitoring requirements and varying metrics quality in microservices QoM Definition Defining Quality of Monitoring (QoM) and implementing an API for cloud provider capabilities in multi-cloud environment
S49 Compute and Network Resource Utilization Monitoring Strengthening surveillance and prediction capabilities of compute and network resource utilization Time Series Resource Utilization Developing algorithms for monitoring and predicting resource utilization of containerized applications using time series data
Category 6: Multi-cloud Coordination and Integration Monitoring
S58 Integrated Multi-Cloud Management Monitoring Implementing integrated monitoring and management across multi-cloud environment Gnocchi Integration Integrating with federated gnocchi for consolidated monitoring and management of multi-cloud environment
S74 Multi-Cloud Monitoring Challenges Addressing monitoring challenges in multi-cloud environment Self-Healing Techniques Implementing a rule-based algorithm for monitoring and self-healing techniques in multi-cloud environment
S84 Cross-Level Monitoring Improvement Ensuring precise and adaptable metrics computation, overcoming inadequate cross-level monitoring approaches Business Process Monitoring Conducting cross-cloud and cross-level business process monitoring and coordinated adaptation of containerized applications
@@ 3.9. Tools and Frameworks (RQ6)
In response to our RQ6 concerning the tools and frameworks utilized for developing containerized multi-cloud applications, an in-depth exploration of the selected studies was conducted. This comprehensive analysis led to the identification of a total of 160 distinct tools and frameworks (see Table 15), all of which are systematically classified into 45 overarching categories, further delineated into 6 subcategories. Notably, the three most prominent categories emerged from this investigation are cloud services and management platform, infrastructure management technologies, and networking tools and framework. Each of these categories encapsulates a range of tools aimed at facilitating the seamless development, deployment, and management of multi-cloud applications. In the following we briefly report the each of the category, shedding light on the tools and frameworks that collectively form the foundation for containerized multi-cloud application development.
Cloud Services and Management Platforms: Cloud Services and Management Platforms: This category of tools facilitates the management and orchestration of cloud services across multi-cloud environments. It includes a broad range of platforms covering Infrastructure as a Service (IaaS), Platform as a Service (PaaS), and Software as a Service (SaaS). These tools are essential for provisioning virtual resources, deploying applications, and delivering software across different cloud platforms. Examples include Amazon EC2, Google Cloud App Engine, and Microsoft Office 365.
Infrastructure Management Technologies: This category of tools facilitates the management of cloud infrastructure. It encompasses container platforms, virtualization solutions, Kubernetes services, orchestration platforms, and tools for configuration management and continuous integration. Key technologies like Docker, Kubernetes, Puppet, Chef, and Terraform play a vital role in automating and optimizing the infrastructure underpinning containerized applications.
Monitoring Tools: This category of tools facilitates the monitoring and analysis of cloud and container operations. It includes solutions for cloud monitoring, container monitoring, visualization & alerting, resource monitoring, and logging & analysis. Tools such as Amazon CloudWatch, cAdvisor, Prometheus, Grafana, and the ELK stack are instrumental in ensuring the performance and reliability of cloud services.
Networking Tools: This category of tools facilitates the establishment and management of network functionalities in a multi-cloud environment. It covers load balancers, reverse proxies, network services, controllers, paradigms, types, and identity protocols, along with storage networks and event buses. Technologies like Software-Defined Networking (SDN), Neutron, and Kafka event bus are crucial for maintaining robust and efficient network operations.
Development and Programming: This category of tools facilitates the development and programming aspects in cloud environments. It includes orchestration tools, various programming languages, integrated development environments (IDEs), service meshes, APIs, and cloud modeling tools. Examples are Java, Python, Eclipse IDE, Istio, REST API, and Cloud Modelling Framework (CLOUDMF), which support the creation and deployment of cloud-native applications.
Security and Assurance: This category of tools facilitates the security and compliance of cloud environments. It encompasses vulnerability management, security platforms & tools, access control & policies, Linux security, and cryptography models & tools. Technologies like OpenVAS Vulnerability Scanner, Kubernetes.eBPF, Seccomp-BPF, and Attribute-Based Encryption and Signature are essential for protecting cloud infrastructure and data against various security threats.
@@
Subcategory Tools and Frameworks
Category 1: Cloud Services and Management Platforms
Cloud Management Platforms
Cloudyn (S65), IBM Brokerage Services (S65), CloudBroker Platform (S69), CloudcheckR (S79), MistIO (S79), ManageIQ (S67, S79), CMP2 (S79), AWS SDK (S71), CloudSME Simulation Platform (CSSP) (S69), MiCADO (S69), WS-PGRADE/gUSE (S69), Vagrant (S08), AWS OpsWorks (S38), Universal Compute Xchange (UCX) (S65), Open vSwitch (OvS) (S72), Liferay portlets (S69), Cloud Service Providers (CSPs) (S71), AWS Data Pipeline (S78), Cloud computing (S25), Amazon AWS (S42, S43), GENI (S38), Google Cloud Platform (S38, S53), Microsoft Azure (S38, S48, S51, S54, S60), AWS EC2 (S51), IBM (S49), Amazon Web Services (S54), Google Cloud (S54), SalesForce (S53), Multi-Cloud (S60), Cloud computing platforms (S74), Multiple independent clouds (S74), Cloud computing (S84)
Cloud Deployment Model
Public clouds: (S37, S53, S59, S61), Google cloud (S07), Microsoft Azure (S08, S16, S19), AWS (S65, S67, S80, S81), Google’s cloud (S65), Google Compute Engine (S67, S80), Alibaba (S65), Tencent (S65), IBM Watson Health Cloud (S78). Private cloud: OpenStack (S20, S21, S43), VMware vSphere Specialized Cloud Platforms: European Grid Infrastructure Federated Cloud (EGI FC) (S81)
Infrastructure as a Service (IaaS)
Amazon EC2 (S01, S51), AWS (S02, S65, S67, S80, S81), Azure (S02, S08, S16, S19, S67, S80), Google GCE (S03, S67, S80), Aliyun ECS (S03), Baidu BCE (S03), Tencent CVM (S03), JD VM (S03), VMware vSphere (S08), GoGrid (S51), Windows Azure (S51)
Platform as a Service (PaaS)
Google Cloud App Engine (S02, S10, S15), Heroku (S08, S53), IBM Bluemix Workflow Service (S17, s53), Amazon Elastic Beanstalk (S20), Cloud Foundry (S24), OpenShift (S24, S53, S55)
Software as a Service (SaaS)
SaaS (S15), Microsoft Office 365 (S53)
Distributed Processing Frameworks
Hadoop (S21, S37), Spark (S21, S28), Apache Yarn (S37), Storm (S21), Apache Kafka (S34, S35)
Data Storage
Apache Cassandra (S28), Firebase real-time database (S35), Redis (S36, S42), MySQL (S42), NoSQL (S85), SQL (S85)
Messaging Systems
Apache Kafka (S34, S35), Zookeeper (S35)
Category 2: Infrastructure Management Technologies
Container Platforms
Docker (S08, S13, S16, S17, S18, S25, S27, S28, S29, S30, S31, S32, S35, S55, S59, S62, S74, S77, S79, S80), Smart Docker Benchmarking Orchestrator (SDBO) (S01), Docklet (S21), LXC (S24, S25, S29), Docker container engine (S24), Docker container (S23), Docker CE version 18.01 (S36), Docker Hub (S36)
Virtualization
Virtual Machines (S05, S22, S23, S46, S72), Amazon EC2 VMs (S18), Docker container virtualization technology (S74), Xen (S19), VMware (S19), Hyper-V (S19), KVM (S19), OpenVZ (S19), Libvirt API (S19)
Kubernetes Services
Microsoft AKS (S03), Aliyun Kubernetes Engine (S03), Tencent Kubernetes Engine (S03)
Orchestration Platforms
Kubernetes (S18, S22, S29, S55, S56, S59, S62, S63, S80), Docker Swarm (S18, S22, S39, S59, S80), Mesos (S21, S22, S59, S80), Yarn (S21), Kubeadm (S33), Kubernetes for container orchestration (S62), Apache Mesos (S59, S80), MiCADO (S69)
Configuration Management
Puppet (S08), Chef (S08, S30), Terraform (S27, S29), Shell scripts (S44), CloudBees (S08) Continuous Integration CloudBees (S08), Maven (S08), NPM (S08), Github (S30), Github (S30)
Container Isolation
Sandboxie (S24), Warden/Garden (S24), Rocket (S24)
Container Infrastructure
Linux Container (S16), Docker Containers on Linux Host (S13), Kubernetes cluster with a Docker (S04), Amazon AWS ECS (S18), Container Technology (S23)
Category 3: Infrastructure Management Technologies
Load Balancer
Load balancer (S23)
Reverse Proxy
Reverse proxy technology (S23), httpd (S36), nginx (S36)
Ports and Interfaces
TrunkPort (S55)
Network Services
Neutron (S55)
Gateways and Tunnels
Gateways (S64), GRE Tunnel (S72)
Network Controllers
Software-Defined Networking (SDN) controller (S72)
Identity Protocols
Host Identity Protocol (HIP) (S77)
Storage Networks
MCSS (S75), Multi-Cloud storage overlay network (S75)
Network Tools
OpenADN (S50), OpenDayLight (S50), Kuryr (S55), Flannel (S55)
Event Buses
Kafka event bus (S34)
Category 4: Development and Programming
Orchestration
MUSA Deployer tool (S52), MUSA Modeler tool (S52), OSM (S55), ManageIQ (S55)
Programming Languages
Java (S30, S34, S35, S54, S80, S81), Go programming language (S32, S36), Python (S35, S54, S56, S58, S79), C# (S54), Ruby (S54), PHP (S54)
IDEs and Editors
Eclipse IDE with tight integration with OPTIMUS (S27), MCSLA editor (S27)
Service Mesh
Consul (S05), Istio (S05), Istio service mesh (S59)
APIs
REST API (S56), OpenStack API (S64)
Cloud Modeling
Cloud Modelling Framework (CLOUDMF) (S20), Cloud Modelling Language (CLOUDML) (S20, S81), CAMEL modeling language (S52, S80, S84)
Category 5: Development and Programming
Vulnerability Management
OpenVAS Vulnerability Scanner (S30)
Security Platforms and Tools
MUSA Security Assurance Platform (S34), Firewall Deployment (S30)
Access Control and Policies
Access control policies (S25)
Linux Security
Linux top command (S34), Kubernetes.eBPF (S40), Seccomp-BPF (S40), Landlock LSM (S40), Pledge (S40), Security Namespace (S40), Linux Security Modules (LSM) (S40)
Cryptography Models and Tools
Attribute-Based Encryption and Signature (S75), ABEABS cryptographic model (S75)
Category 6: Monitoring Tools
Cloud Monitoring
Amazon CloudWatch (S19)
Container Monitoring
cAdvisor (S22), Prometheus (S22), AlertManager (S22)
Visualization and Alerting
Grafana (S27), Telegraf (S27)
Resource Monitoring
Resource monitoring tool (S60), Ceilometer (S56, S58), Gnocchi (S58)
Logging and Analysis
ELK stack (S74)

### v2

@@ 3.9. Tools and Frameworks (RQ6)
In response to RQ6 regarding tools and frameworks for developing containerized multi-cloud applications, we identified 87 distinct tools and frameworks, classified into 5 subcategories (see Table 15 and the Tools and Frameworks sheet in (Waseem et al., 2024)). Below, we provide an overview of each category, along with key tools and frameworks, highlighting their functionality, adoption, and distinctive features.
Container Orchestration and Management Tools: are essential for automating the deployment, scaling, and management of containerized applications across distributed and multi-cloud environment. These tools assist developers in managing large-scale container deployments and efficiently allocating resources across different cloud providers. Among the most popular orchestration and management tools, Kubernetes is widely adopted for orchestrating containers at scale, providing high availability, and facilitating load balancing across nodes. Docker and its variants, such as Docker Swarm, offer more convenient container orchestration but are less scalable than Kubernetes. Mesos is another orchestration tool that supports not only containers but also other workloads, such as big data processing tasks. Tools like LXC and runC provide the runtime environment to run a single container with fine-grain control. Overall, this category covers key orchestration and container management tools that address diverse deployment needs in multi-cloud environment.
Multi-cloud and Hybrid Cloud Platforms: are designed to support organizations in deploying and managing applications across different cloud providers smoothly. In this category, we identify and organize tools like Cloudify and CMP2, which offer end-to-end multi-cloud orchestration capabilities, along with platforms such as Cloudcheckr and MistIO, which provide monitoring and cost management across various cloud services. Similarly, AWS SDK and the CloudSME Simulation Platform (CSSP) are widely adopted for interacting with cloud APIs and simulating cloud environments, respectively. Moreover, hybrid platforms like OpenStack and VMware vSphere offer a blend of public and private cloud deployment options, which are essential for organizations pursuing a hybrid cloud strategy. This category highlights the tools that enable integration, scalability, and cost management for applications running across multiple clouds.
Networking, Security, and Access Management Tools: This category represents the tools that help to minimize security risks and ensure connectivity across containerized workloads in multi-cloud systems. As multi-cloud deployments become increasingly complex, advanced networking tools like Open vSwitch (OvS) enable developers to manage virtual networks across various cloud platforms. Security tools such as Seccomp-BPF and Linux Security Modules (LSM) provide critical protection to containers by enforcing strict access controls and sandboxing. Identity management solutions, such as Host Identity Protocol (HIP) and access control policies, ensure proper authentication and authorization for users and services in the cloud.
Monitoring, Management, and Automation Tools: Tools in this category provide the foundation for a scalable and automated approach to managing distributed applications by ensuring that workloads are continuously optimized and monitored on cloud platforms. Tools like Prometheus and Grafana offer powerful visualization capabilities for monitoring CPU usage, memory consumption, and network activity, aiding system administrators. With Terraform and Puppet, it is possible to automate infrastructure management, enabling developers to define infrastructure as code and automate provisioning across different cloud environments. Other tools, such as Ceilometer and Gnocchi, are responsible for gathering telemetry data in large OpenStack deployments.
Development, APIs, and Cloud Storage Solutions: This category organizes the main tools that support software development, API integration, and data storage in containerized multi-cloud environment. The tools include REST API and YAML, which enable services and applications deployed across multiple cloud providers to communicate for data sharing and inter-service interactions. General programming languages such as Java, Go, and Python are commonly used to develop applications based on cloud-native architecture, alongside tools for version control, such as GitHub, and project management using Maven. On the data storage side, tools like gcloud storage and azblob provide scalable and fault-tolerant systems for managing large datasets in cloud environments.
@@
Category Tool, Framework, and Language
Container Orchestration and Management Kubernetes (Pai and Kumar, 2021; Wong et al., 2019; Gupta et al., 2021; Bolivar et al., 2018; del Castillo et al., 2013; Rodigari et al., 2021; Rahman and Lama, 2019; Pellegrini et al., 2017; Quint and Kratzke, 2018; Barletta et al., 2024; Struhár et al., 2024; Barati et al., 2023; Horchulhack et al., 2024; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Pandi et al., 2023; Sofia et al., 2024; Lee and Nam, 2023; Sofia et al., 2023; Nguyen and Chien, 2023; Tiwari and Sharma, 2023; Santos et al., 2023; Roy and Kalafatis, 2023; Kataru et al., 2023; Benmerar et al., 2023; Alonso et al., 2023a; Harzenetter et al., 2023; Huang and Wang, 2023; Wang et al., 2023; Nagarajan et al., 2023), Docker (Pham et al., 2015; Naik, 2016a; Ouyang et al., 2022; Slominski et al., 2015; Pai and Kumar, 2021; Gao et al., 2017; Alonso et al., 2019; Barna et al., 2017; Gupta et al., 2021; Sheridan et al., 2017; Mittal and Risco-Martín, 2017; Al-shammari and Alwan, 2018; Carpio et al., 2020; Bolivar et al., 2018; Rodigari et al., 2021; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ranjbar et al., 2017; Serhiienko and Spillner, 2018; Quint and Kratzke, 2018), Docker Swarm (Pai and Kumar, 2021; Wong et al., 2019; Altran et al., 2022; Rodigari et al., 2021; Quint and Kratzke, 2018), Mesos (Cui et al., 2016; Wong et al., 2019; Rodigari et al., 2021; Quint and Kratzke, 2018), LXC (Pahl, 2015; Gao et al., 2017; Gupta et al., 2021), OverlayFS (Al-shammari and Alwan, 2018), runC (Bélair et al., 2019), Container Technology (Wang et al., 2020), Docklet (Cui et al., 2016)
Multi-cloud and Hybrid Cloud Platforms multi-cloud environment (Shi et al., 2020; Fadda et al., 2019; Di Modica et al., 2019; Tang, 2021), Cloudify (Zeck and Bouroudjian, 2017), CloudBroker Platform (Kiss, 2018), Cloudcheckr (Serhiienko and Spillner, 2018), MistIO (Serhiienko and Spillner, 2018), ManageIQ (Tang, 2021; Serhiienko and Spillner, 2018), CMP2 (Serhiienko and Spillner, 2018), AWS SDK (Graupner et al., 2015a), CloudSME Simulation Platform (CSSP) (Kiss, 2018), Vagrant (Pham et al., 2015), AWS OpsWorks (Pham et al., 2015), Universal Compute Xchange (UCX) (Zeck and Bouroudjian, 2017), Liferay portlets (Kiss, 2018), Cloud Service Providers (CSPs) (Graupner et al., 2015a), AWS Data Pipeline (Sitaram et al., 2018), GENI (Pandey et al., 2022), Google Cloud Platform (Pandey et al., 2022; Acquaviva et al., 2017; Tang, 2021; Quint and Kratzke, 2018), AWS EC2 (He et al., 2022; Buyya and Barreto, 2015; Pai and Kumar, 2021), Alibaba cloud (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), Tencent (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), OpenStack (Cui et al., 2016; Dreibholz et al., 2019), VMware vSphere (Mfula and Nurminen, 2018), Aliyun ECS (Hua et al., 2020), Baidu BCE (Hua et al., 2020), Tencent CW (Hua et al., 2020), JD VW (Hua et al., 2020), GoGrid (Buyya and Barreto, 2015)
Networking, Security, and Access Management Open vSwitch (OvS) (Alaluna et al., 2017), Cloud networks (Ranjbar et al., 2017), Host Identity Protocol (HIP) (Ranjbar et al., 2017), Gateways (Harwalkar et al., 2019a), Pledge (Bélair et al., 2019), Seccomp-BPF (Bélair et al., 2019), Landlock LSM (Bélair et al., 2019), Linux Security Modules (LSM) (Bélair et al., 2019), MUSA Security Assurance Platform (Rios et al., 2017), Firewall Deployment (Sheridan et al., 2017), Access control policies (Gao et al., 2017), Attribute-Based Encryption and Signature (Shen et al., 2017), ABE/ABS cryptographic model (Shen et al., 2017)
Monitoring, Management, and Automation Prometheus (Wong et al., 2019; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Grafana (Alonso et al., 2019; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Telegraf (Alonso et al., 2019), AWS CloudWatch (Alhamazani et al., 2014), cAdvisor (Wong et al., 2019), Ceilometer (del Castillo et al., 2013; Sitaram et al., 2018), Gnocchi (Sitaram et al., 2018), ELK stack (Mfula and Nurminen, 2018), Terraform (Alonso et al., 2019; Gupta et al., 2021; da Silva et al., 2023; Pandi et al., 2023; Harzenetter et al., 2023), CloudBees (Pham et al., 2015), Jenkins (Pandey et al., 2022; Solayman and Qasha, 2023), Puppet (Pham et al., 2015), Chef (Pham et al., 2015; Sheridan et al., 2017), GitHub (Sheridan et al., 2017), Shell scripts (Rios et al., 2017), Maven (Pham et al., 2015), NPM (Pham et al., 2015), Consul (Pham et al., 2015)
Development, APIs, and Cloud Storage Public cloud storage (Harwalkar et al., 2019a), Private cloud storage (Harwalkar et al., 2019a), Hybrid cloud storage (Harwalkar et al., 2019a), Google Cloud Storage (Ramesh et al., 2023), Azure Blob Storage (Ramesh et al., 2023), REST API (Harwalkar et al., 2019a), YAML (Serhiienko and Spillner, 2018), OpenVAS Vulnerability Scanner (Sheridan et al., 2017), DAX format (Tang, 2021), Java (Sheridan et al., 2017; Rios et al., 2017; Carpio et al., 2020; Bucur et al., 2018; Quint and Kratzke, 2018; Challita et al., 2017), Go programming language (Al-shammari and Alwan, 2018; Lorenzo et al., 2019), Python (Carpio et al., 2020; Bucur et al., 2018; del Castillo et al., 2013; Sitaram et al., 2018; Serhiienko and Spillner, 2018), Ruby (Bucur et al., 2018), PHP (Bucur et al., 2018), C# (Bucur et al., 2018), Cloud Modelling Framework (CLOUMDF) (Ferry et al., 2014), Cloud Modelling Language (CLOUDML) (Ferry et al., 2014; Challita et al., 2017), CAMEL modeling language (Casola et al., 2017; Quint and Kratzke, 2018; Kritikos et al., 2019), Eclipse IDE (Alonso et al., 2019), MCSLA editor (Alonso et al., 2019)